OpenStudy (anonymous):

last question for tonight Let f(x) = 3 x^2 - 11. Find an equation of the line tangent to the graph of f at the point where x =-2 y=

OpenStudy (anonymous):

i know i find the slope

sam (.sam.):

dy/dx = 6x x=-2, dy/dx= -12

sam (.sam.):

y-1=-12(x-(-2)) y=-12x-23

sam (.sam.):

Anything not sure just ask.

OpenStudy (anonymous):

ok then thanks im studying problems in my online hw for a test friday

OpenStudy (anonymous):

f(x) = 3 x^2 - 11. f( -2) = 3 * 4 - 11 = 1 f' (x) = 6x slope m = f'( -2) = 6(-2) = -12 => Tangent line y -1 = -12 ( x +2) y = -12x -23